CoronaV Corp. just paid $3 dividends, it was $2.15 four years ago. The dividends should continue to grow at the same rate to infinity. How much is their stock price given that investors require 12% returns? (Hint: Use 2 decimals for the dividend growth rate.)

Question 1 options:

$34.69

$53.16

$128.33

$98.51
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

3=2.15*(1+r/100)^4

(3/2.15)^(1/4)=(1+r/100)

(1+r/100)=1.0869

r=1.0869-1

=8.69%(Approx)

Current price=D1/(Required return-Growth rate)

=(3*(1+0.0869))/(0.12-0.0869)

=$98.51(Approx).